bryan berg built an 83 story house of cards in 1995 he used 27000 playing cards to accomplish this feat. At this same rate about how many cards would it take to build a 100 story house of cards

Answer by Johngift(2)

1 storey=27000/83=325.3 cards.100 storey=325.3*100=32530.he needs 32530 cards.